OpenStudy (angel123):

The Leafs’ coach stated that “the likelihood of us winning the next game are 2/9, the likelihood of tying the next game are 1/2, and likelihood of us loosing the next game are 2/5”. Can the coach’s predictions be entirely correct? Justify your response

OpenStudy (vuriffy):

Common denominator to see if the coach's predictions were corrected.

OpenStudy (vuriffy):

Lowest value in which 9, 5, 2; all go in one number?

OpenStudy (vuriffy):

Any guess?

OpenStudy (angel123):

i m not sure i dont get the question but i guess what ur saying could work

OpenStudy (vuriffy):

Use this little technique, in any case, you could multiply all three. 9*5*2 = 90. yes?

OpenStudy (vuriffy):

Understand?

OpenStudy (vuriffy):

All 3 go into 90.

OpenStudy (angel123):

yea thank you!!

OpenStudy (vuriffy):

Now to further elaborate, you must multiply the numerator by the value that number is able to go into 90.

OpenStudy (vuriffy):

So, 90/9 = 10, 90/5 = 18, 90/2 = 45.

OpenStudy (vuriffy):

Apologies, my wifi suddenly dropped.

OpenStudy (angel123):

no you helped alot thanks alot

OpenStudy (vuriffy):

Now toward the original numerators you gave me: Multiply each top number by the correlating values which I have given. Can you calculate that out for me please?

OpenStudy (angel123):

so should i do 10+18+45=73 and then divide it by 90?

OpenStudy (vuriffy):

No Angel. Let me clear it up a little more?

OpenStudy (angel123):

no wait i got it give me one sec

OpenStudy (vuriffy):

2/9, 1/2, 2/5 Denominators are now in /90 10*2 = ? 1*45 = ? 2*18 = ? Then add those all up for the numerator.

OpenStudy (angel123):

yea so silly of me so it would be 20+45+36=101

OpenStudy (vuriffy):

Yes, correct. 101/90 does not equal 1. If it was 90/90, then the coach would be correct. It must equal one in this type of question.

OpenStudy (angel123):

oh okay i got it now thanks alot once again

OpenStudy (vuriffy):

Of course, anytime!

OpenStudy (vuriffy):

@Nnesha Am I pro at OpenStudy yet?

Nnesha (nnesha):

you're pro at planning& gaming.